# reference.bib

### v2

@@ V-A Open Challenges
Although significant research has been conducted by researchers in the field of AD, and the AD industry is flourishing, there are still open challenges that require researchers’ attention to achieve fully intelligent autonomous driving. These challenges have been individually discussed with support from relevant literature:
@@ V-A1 Coarse-Structured Information
Coarse-Structured Information: Most of the datasets introduced in the literature for 3D lane detection in AD are recorded in normal and well-structured infrastructures of advanced cities. The currently developed deep learning models may achieve the best results over structured datasets, but they generalize poorly in many unstructured environments. This issue of AD demands further attention in terms of data collection, as well as the inclusion of new and effective representation mechanisms in deep learning models.
@@ V-A2 Uncertainty-Aware Decisions
A largely overlooked aspect in lane detection and AD decision-making is the confidence with which models make predictions on input data. However, the confidence of model outputs plays a crucial role in ensuring the safety of autonomous driving. The inherently uncertain nature of the vehicular surroundings seems to have not convinced the community to delve into this matter, as the current methodological trends focus solely on predictive scores. Fortunately, confidence estimation has recently gained attention in the community [arnez2020comparison, michelmore2018evaluating]. Nevertheless, elements from evidential deep learning [sensoy2018evidential], Bayesian formulations of deep neural networks [kendall2017uncertainties], simpler mechanisms to approximate the output confidence of neural networks (e.g., Monte Carlo dropout [gal2016dropout] or ensembles [lakshminarayanan2017simple]), and other assorted methods for uncertainty quantification [abdar2021review] should be progressively incorporated as an additional yet crucial criterion for decision making. When dealing with complex environments, where there is a lack of data that fully represents all possible scenes, the model outputs a significant amount of epistemic uncertainty. Without considering confidence as an additional factor for AD or with current studies focused solely on predictive and/or computational efficiency aspects, there is no guarantee that the emerging 3D lane detection models in the scientific community will be practically useful and transferable to the industry.
@@ V-A3 Weakly supervised Learning Strategies
In current deep learning-based models, the majority rely on fully supervised learning strategies, which have high demands for labeled data. This is particularly challenging in the field of 3D lane detection, as general visual sensor data lacks depth information. It is difficult to simply assign 3D information to lanes based on images alone, and alternative sensors such as LiDAR are required to obtain the 3D lane information. This leads to the costly and labor-intensive task of annotating 3D lane data. Fortunately, both the academic and industrial communities have recognized this issue, and weakly supervised learning strategies have received extensive research and attention in the field of deep learning. However, in the specific branch of 3D lane detection, there is currently limited research focusing on weakly supervised learning strategies. If we can effectively leverage self-supervised/weakly supervised learning strategies, it would significantly reduce data collection costs and allow for more training data to enhance the performance of 3D lane detection algorithms, thereby further advancing the autonomous driving industry.
@@ V-B1 Video-Based 3D Lane Detection for Autonomous Driving
Drawing from the advancements in video-based object detection [wu2019sequence], semantic segmentation [wang2021swiftnet, maninis2018video, hu2018videomatch], and 2D lane detection [Jin2023RecursiveVL, zhang2021vil, zhang2018end], it’s evident that incorporating video-based techniques significantly improves the precision and dependability of 3D lane detection systems. The core advantage of video-based methods is their capacity to harness temporal data, providing a dynamic view that static images lack. This dynamic perspective is particularly vital in comprehending and predicting complex driving situations in three-dimensional spaces, where the intricacies of lane placement and vehicular interactions intensify. Methods like Recursive Video Lane Detection (RVLD) [Jin2023RecursiveVL] demonstrate the power of video in capturing ongoing lane shifts and variations over time, a feature immensely beneficial for 3D modeling accuracy. Additionally, incorporating video data into these systems enhances our understanding of spatial dynamics in driving environments, which is crucial for 3D lane detection. By incorporating sophisticated deep learning techniques used in video-based object detection and semantic segmentation, future iterations of 3D lane detection systems could attain advanced spatial awareness, significantly enhancing the navigation capabilities and safety of autonomous vehicles.
@@ V-B2 Hybrid Approaches and Multimodality
The progression of multi-modal 3D lane detection technologies for autonomous vehicles is greatly accelerated by the incorporation of various sensory inputs, such as cameras, LiDAR, and radar. This integration marks a promising avenue for overcoming the challenges faced by existing camera-reliant systems. This approach, underscored by the successes in multi-modal 3D object detection and semantic segmentation [feng2020deep], leverages the complementary strengths of each sensor type to enhance detection accuracy and reliability, particularly in challenging environmental and complex driving scenarios [chen2019progressive, liang2019multi]. Reflecting on pioneering models such as “Deep Multi-Sensor Lane Detection” [bai2018deep] and “$M^{2}$-3DLaneNet” [luo2022m], which have effectively utilized multi-sensor inputs to refine lane boundary estimations and demonstrated robust performance against occlusions and variable lighting conditions, the potential for significant advancements in lane detection technology is evident. The future trajectory in this field should emphasize the exploration of advanced data fusion methods, meticulous sensor calibration, and synchronization techniques, alongside the exploitation of emerging technologies like edge computing for real-time multi-modal data processing.
@@ V-B3 Active Learning and Incremental Learning
Active learning [ren2021survey] in machine learning refers to the ability of a model to adapt and learn over time and encountering new data during the testing phase and after deployment. In real-world environments, vehicles may encounter randomly occurring unfamiliar scenes and lane topology, which may require AI models to make decisions for further actions such as braking or accelerating to achieve sensible driving operations. Therefore, lane detection techniques should allow for an interactive approach to handle various types of scenes and lane topologies, involving human annotators to label unlabeled data instances and human involvement in the training process [khan2022pmal]. There are different types of active learning techniques, such as membership query synthesis [kelner2020learning], where synthetic data is generated and the parameters of the synthetic data can be adjusted based on the structure of the data [ahmed2022aaqal], which originates from the underlying species of the dataset. On the other hand, the ability of 3D lane detection models to incrementally update their captured knowledge with new data is crucial for their sustainability and continuous improvement. We anticipate that both these capabilities of 3D lane detection models in road understanding will become increasingly important in future research.
@@ V-B4 Adverse Weather Conditions
The progress in developing camera-based 3D lane detection systems for autonomous driving is notably hindered by adverse weather conditions, which significantly impair visibility. Events like heavy rainfall, fog, snow, and dust storms can severely impact the functionality of these systems [zhang2023perception]. The core issue stems from the compromised quality of visual data necessary for the precise detection and segmentation of lane markings, resulting in a decline in reliability and an increase in the likelihood of false negatives or positives. This diminishment in the efficacy of such systems not only elevates safety hazards but also constrains the operational scope of autonomous vehicles. However, recent breakthroughs in object detection and segmentation, as illustrated in research works such as “ACDC: The Adverse Conditions Dataset with Correspondences for Semantic Driving Scene Understanding” [sakaridis2021acdc] and “Vehicle Detection and Tracking in Adverse Weather Using a Deep Learning Framework” [9184996], reveal avenues for enhancing 3D lane detection amidst challenging weather conditions. These studies put forth the advantages of employing deep learning algorithms trained on datasets that include a wide array of adverse weather instances, showcasing the significance of effective data augmentation, domain adaptation tailored to specific conditions, and the employment of semantic segmentation techniques. Employing these approaches, the capabilities of camera-based detection systems can be substantially improved to interpret lane markings accurately and ensure safe navigation even under poor visibility, laying down an optimistic pathway for ongoing research and development in the field of autonomous driving technology.
@@
Figure 9: The test results of the current open-source state-of-the-art LATR [luo2023latr] algorithm under extreme weather conditions or photography scenarios are depicted in the figure above, with red lines indicating the algorithm’s predicted lane lines and blue lines representing the ground truth lane lines. Figures a and d show scenes with strong exposure during the day and at night, respectively; figures b and e depict rainy scenes during the day and night; and figures c and f show foggy scenes during the day and night. From these test scenarios under harsh conditions, it is evident that the current algorithm performs poorly in detecting lane lines, indicating that there is substantial room for improvement.Additionally, there is a lack of datasets for training and testing under such severe conditions.
V-B5 3D Lane Detection with Large Language Model (LMM)
The advent of large language models (LLMs) such as ChatGPT [wu2023brief] has transformed the field of artificial general intelligence (AGI), demonstrating their remarkable zero-shot abilities in tackling a variety of natural language processing (NLP) tasks using customized user prompts or language instructions.
@@
Computer vision encompasses a set of unique challenges and concepts that stand apart from those found in NLP. Vision foundation models typically adhere to a pre-training and subsequent fine-tuning process [wang2023internimage, chen2022vision, su2023towards, wang2022image, fang2023eva, tao2023siamese], which, while effective, entails significant additional costs for adaptation to a range of downstream applications. Techniques such as multi-task unification [radford2018improving, wang2022ofa, alayrac2022flamingo, wang2022git] aim to endow systems with a broad set of capabilities, yet they often fail to transcend the constraints of predetermined tasks, leaving a noticeable capability deficit in open-ended tasks when juxtaposed with LLMs. The advent of visual prompt tuning [jia2022visual, yao2021cpt, zang2022unified, wang2023images] provides a novel method for delineating specific vision tasks, like object detection, instance segmentation, and pose estimation, through the use of visual masking. Nonetheless, there is currently no work that combines LLMs with 3D lane line detection. As large language models become more widespread and their capabilities continue to improve, research into LLM-based lane line detection presents an interesting and promising avenue for future exploration.
Balanced Scene 3DLaneNet [garnett20193d] 53 RTX 2080 89.3 86.4 0.068 0.477 0.015 0.202
Gen-LaneNet [guo2020gen] 60 RTX 2080 90.1 88.1 0.061 0.496 0.012 0.214
CLGo [liu2022learning] 75 RTX 2080 94.2 91.9 0.061 0.361 0.029 0.250
PersFormer [chen2022persformer] 21 Tesla V100 - 92.9 0.054 0.356 0.010 0.234
Anchor3DLane [huang2023anchor3dlane] 87.3 RTX 2080Ti 97.2 95.6 0.052 0.306 0.015 0.223
BEV-LaneDet [wang2023bevlanedet] 102 Tesla V100 - 98.7 0.016 0.242 0.020 0.216
LATR [luo2023latr] 17.75 Tesla A100 97.9 96.8 0.022 0.253 0.007 0.202
D-3DLD [kim2023d] 22.6 Tesla A100 95.5 93.3 0.039 0.412 0.011 0.234
Rarely Observed 3DLaneNet [garnett20193d] 53 RTX 2080 74.6 72.0 0.166 0.855 0.039 0.521
Gen-LaneNet [guo2020gen] 60 RTX 2080 79.0 78.0 0.139 0.903 0.030 0.539
CLGo [liu2022learning] 75 RTX 2080 88.3 86.1 0.147 0.735 0.071 0.609
PersFormer [chen2022persformer] 21 Tesla V100 - 87.5 0.107 0.782 0.024 0.602
Anchor3DLane [huang2023anchor3dlane] 87.3 RTX 2080Ti 96.9 94.4 0.094 0.693 0.027 0.579
BEV-LaneDet [wang2023bevlanedet] 102 Tesla V100 - 99.1 0.031 0.594 0.040 0.556
LATR [luo2023latr] 17.75 Tesla A100 97.3 96.1 0.050 0.600 0.015 0.532
Visual Variations 3D-LaneNet [garnett20193d] 53 RTX 2080 74.9 72.5 0.115 0.601 0.032 0.230
Gen-LaneNet [guo2020gen] 60 RTX 2080 87.2 85.3 0.074 0.538 0.015 0.232
CLGo [liu2022learning] 75 RTX 2080 89.2 87.3 0.084 0.464 0.045 0.312
PersFormer [chen2022persformer] 21 Tesla V100 - 89.6 0.074 0.430 0.015 0.266
Anchor3DLane [huang2023anchor3dlane] 87.3 RTX 2080Ti 93.6 91.4 0.068 0.367 0.020 0.232
BEV-LaneDet [wang2023bevlanedet] 102 Tesla V100 - 96.9 0.027 0.320 0.031 0.256
LATR [luo2023latr] 17.75 Tesla A100 96.6 95.1 0.045 0.315 0.016 0.228
TABLE IV: An in-depth comparison of various methods on the ApolloSim dataset under three different split settings. “N” and “F” are short for near and far respectively.
3D-LaneNet [garnett20193d] 53 RTX 2080 44.1 40.8 46.5 47.5 41.5 32.1 41.7
GenLaneNet [guo2020gen] 50 RTX 2080 32.3 25.4 33.5 28.1 18.7 21.4 31.0
PersFormer [chen2022persformer] 21 Tesla V100 50.5 42.4 55.6 48.6 46.6 40.0 50.7
Anchor3DLane [huang2023anchor3dlane] 87.3 RTX 2080Ti 54.3 47.2 58.0 52.7 48.7 45.8 51.7
BEV-LaneDet [wang2023bevlanedet] 102 Tesla V100 58.4 48.7 63.1 53.4 53.4 50.3 53.7
GroupLane [li2023grouplane] 38.2 RTX 2080 64.1 - - - - - -
LATR [luo2023latr] 17.75 Tesla A100 61.9 55.2 68.2 57.1 55.4 52.3 61.5
DecoupleLane [han2023decoupling] - - 51.2 43.5 57.3 - 48.9 43.5 -
TABLE V: Comparison of various 3D methods on the OpenLane benchmark, evaluating the F-Score on both the entire validation set and separate sets corresponding to five different scenarios.
Method FPS Device F1(%) Precision(%) Recall(%) CD Error(m)
3D-LaneNet [garnett20193d] 53 RTX 2080 44.73 61.46 35.16 0.127
Gen-LaneNet [guo2020gen] 60 RTX 2080 45.59 63.95 35.42 0.121
SALAD [yan2022once] - - 64.07 75.90 55.42 0.098
PersFormer [chen2022persformer] 21 Tesla V100 74.33 80.30 69.18 0.074
WS-3D-Lane [ai2022ws] 26.7 RTX 3080Ti 77.02 84.51 70.75 0.058
Anchor3DLane [huang2023anchor3dlane] 87.3 RTX 2080Ti 74.87 80.85 69.71 0.060
GroupLane [li2023grouplane] 38.2 RTX 2080 80.73 82.56 78.90 0.053
LATR [luo2023latr] 17.75 Tesla A100 80.59 86.12 75.73 0.052
DecoupleLane [han2023decoupling] - - 75.07 81.19 69.26 0.062
TABLE VI: The performance of 3D lane detection on the ONCE-3DLane dataset. CD error represents the Chamfer distance between the predicted lane line and the ground truth.
@@ V-B6 Towards More Accurate and Efficient 3D Lane Detection Methods for Autonomous Driving
The qualitative performance of current 3D lane detection techniques is shown in Table IV. We can observe that only a few methods manage to strike a balance between model accuracy and inference latency. The experimental results of these methods indicate the need for further improvements to alleviate computational burdens while maintaining their unparalleled performance. Additionally, we selected some challenging data from the 3D lane detection dataset and tested the performance of the 3D lane line detection algorithm on these challenging data samples. However, the performance of the algorithm was not satisfactory,as depicted in Fig. 9. Improving the algorithm’s detection performance in extreme weather conditions is also crucial. Moreover, the reported time complexities in Table IV,V and VI suggest that some methods can achieve real-time execution when deployed on GPU devices. Nevertheless, considering the constrained computational resources in today’s autonomous driving system, the focus of 3D lane detection methods should also shift towards computational complexity.

### v3

@@ V-A Open Challenges
Although significant research has been conducted by researchers in the field of AD, and the AD industry is flourishing, there are still open challenges that require researchers’ attention to achieve fully intelligent AD. These challenges have been individually discussed with support from relevant literature:
@@ V-A1 Coarse-Structured Information
Most of the datasets introduced in the literature for 3D lane detection in AD are recorded in normal and well-structured infrastructures of advanced cities. The currently developed deep learning models may achieve the best results over structured datasets, but they generalize poorly in many unstructured environments. This issue of AD demands further attention in terms of data collection, as well as the inclusion of new and effective representation mechanisms in deep learning models.
@@ V-A2 Uncertainty-Aware Decisions
A largely overlooked aspect in lane detection and AD decision-making is the confidence with which models make predictions on input data. However, the confidence of model outputs plays a crucial role in ensuring the safety of AD. The inherently uncertain nature of the vehicular surroundings seems to have not convinced the community to delve into this matter, as the current methodological trends focus solely on predictive scores. Fortunately, confidence estimation has recently gained attention in the community [arnez2020comparison, michelmore2018evaluating]. Nevertheless, elements from evidential deep learning [sensoy2018evidential], Bayesian formulations of deep neural networks [kendall2017uncertainties], simpler mechanisms to approximate the output confidence of neural networks (e.g., Monte Carlo dropout [gal2016dropout] or ensembles [lakshminarayanan2017simple]), and other assorted methods for uncertainty quantification [abdar2021review] should be progressively incorporated as an additional yet crucial criterion for decision making. When dealing with complex environments, where there is a lack of data that fully represents all possible scenes, the model outputs a significant amount of epistemic uncertainty. Without considering confidence as an additional factor for AD or with current studies focused solely on predictive and/or computational efficiency aspects, there is no guarantee that the emerging 3D lane detection models in the scientific community will be practically useful and transferable to the industry.
@@ V-A3 Weakly Supervised Learning Strategies
In current deep learning-based models, the majority rely on fully supervised learning strategies, which have high demands for labeled data. This is particularly challenging in the field of 3D lane detection, as general visual sensor data lacks depth information. It is difficult to simply assign 3D information to lanes based on images alone, and alternative sensors such as LiDAR are required to obtain the 3D lane information. This leads to the costly and labor-intensive task of annotating 3D lane data. Fortunately, both the academic and industrial communities have recognized this issue, and weakly supervised learning strategies have received extensive research and attention in the field of deep learning. However, in the specific branch of 3D lane detection, there is currently limited research focusing on weakly supervised learning strategies. If we can effectively leverage self-supervised/weakly supervised learning strategies, it would significantly reduce data collection costs and allow for more training data to enhance the performance of 3D lane detection algorithms, thereby further advancing the AD industry.
@@ V-B1 Video-Based 3D Lane Detection for Autonomous Driving
Drawing from the advancements in video-based object detection [wu2019sequence], semantic segmentation [wang2021swiftnet, maninis2018video, hu2018videomatch], and 2D lane detection [Jin2023RecursiveVL, zhang2021vil, zhang2018end], it is evident that incorporating video-based techniques significantly improves the precision and dependability of 3D lane detection systems. The core advantage of video-based methods is their capacity to harness temporal data, providing a dynamic view that static images lack. This dynamic perspective is particularly vital in comprehending and predicting complex driving situations in 3D spaces, where the intricacies of lane placement and vehicular interactions intensify. Methods like Recursive Video Lane Detection (RVLD) [Jin2023RecursiveVL] demonstrate the power of video in capturing ongoing lane shifts and variations over time, a feature immensely beneficial for 3D modeling accuracy. Additionally, incorporating video data into these systems enhances our understanding of spatial dynamics in driving environments, which is crucial for 3D lane detection. By incorporating sophisticated deep learning techniques used in video-based object detection and semantic segmentation, future iterations of 3D lane detection systems could attain advanced spatial awareness, significantly enhancing the navigation capabilities and safety of autonomous vehicles.
@@ V-B2 Hybrid Approaches and Multimodality
The progression of multi-modal 3D lane detection technologies for autonomous vehicles is greatly accelerated by the incorporation of various sensory inputs, such as cameras, LiDAR, and radar. This integration marks a promising avenue for overcoming the challenges faced by existing camera-reliant systems. This approach, underscored by the successes in multi-modal 3D object detection and semantic segmentation [feng2020deep], leverages the complementary strengths of each sensor type to enhance detection accuracy and reliability, particularly in challenging environmental and complex driving scenarios [chen2019progressive, liang2019multi]. Reflecting on pioneering models such as [bai2018deep] and [luo2022m], which have effectively utilized multi-sensor inputs to refine lane boundary estimations and demonstrated robust performance against occlusions and variable lighting conditions, the potential for significant advancements in lane detection technology is evident. The future trajectory in this field should emphasize the exploration of advanced data fusion methods, meticulous sensor calibration, and synchronization techniques, alongside the exploitation of emerging technologies like edge computing for real-time multi-modal data processing.
@@ V-B3 Active Learning and Incremental Learning
Active learning [ren2021survey] in machine learning allows a model to adapt and learn over time. This learning process takes place when the model encounters new data during testing and after deployment, making it especially valuable in real-world environments. In real-world environments, vehicles may encounter randomly occurring unfamiliar scenes and lane topology, which may require artificial intelligence models to make decisions for further actions such as braking or accelerating to achieve sensible driving operations. Therefore, lane detection techniques should allow for an interactive approach to handle various types of scenes and lane topologies, involving human annotators to label unlabeled data instances and human involvement in the training process [khan2022pmal]. There are different types of active learning techniques, such as membership query synthesis [kelner2020learning], where synthetic data is generated and the parameters of the synthetic data can be adjusted based on the structure of the data [ahmed2022aaqal], which originates from the underlying species of the dataset. On the other hand, the ability of 3D lane detection models to incrementally update their captured knowledge with new data is crucial for their sustainability and continuous improvement. We anticipate that both these capabilities of 3D lane detection models in road understanding will become increasingly important in future research.
@@ V-B4 Adverse Weather Conditions
The progress in developing camera-based 3D lane detection systems for AD is notably hindered by adverse weather conditions, which significantly impair visibility. Events like heavy rainfall, fog, snow, and dust storms can severely impact the functionality of these systems [zhang2023perception]. The core issue stems from the compromised quality of visual data necessary for the precise detection and segmentation of lane markings, resulting in a decline in reliability and an increase in the likelihood of false negatives or positives. This diminishment in the efficacy of such systems not only elevates safety hazards but also constrains the operational scope of autonomous vehicles. However, recent breakthroughs in object detection and segmentation, as illustrated in research works such as [sakaridis2021acdc] and [9184996], reveal avenues for enhancing 3D lane detection amidst challenging weather conditions. These studies put forth the advantages of employing deep learning algorithms trained on datasets that include a wide array of adverse weather instances, showcasing the significance of effective data augmentation, domain adaptation tailored to specific conditions, and the employment of semantic segmentation techniques. Employing these approaches, the capabilities of camera-based detection systems can be substantially improved to interpret lane markings accurately and ensure safe navigation even under poor visibility, laying down an optimistic pathway for ongoing research and development in the field of AD technology.
@@
Figure 10: The test results of the current open-source state-of-the-art LATR [luo2023latr] algorithm under extreme weather conditions or photography scenarios are depicted, with red lines indicating the algorithm’s predicted lane lines and blue lines representing the ground truth lane lines. (a) and (d) show scenes with strong exposure during the day and at night, respectively; figures (b) and (e) depict rainy scenes during the day and night; and figures (c) and (f) show foggy scenes during the day and night. From these test scenarios under harsh conditions, it is evident that the current algorithm performs poorly in detecting lane lines, indicating that there is substantial room for improvement. Additionally, there is a lack of datasets for training and testing under such severe conditions.
V-B5 3D Lane Detection with Large Language Model
@@
The advent of (LLMs), such as ChatGPT [wu2023brief], have transformed the field of artificial general intelligence (AGI), demonstrating their remarkable zero-shot abilities in tackling a variety of natural language processing (NLP) tasks using customized user prompts or language instructions. Computer vision encompasses a set of unique challenges and concepts that stand apart from those found in NLP. Vision foundation models typically adhere to a pre-training and subsequent fine-tuning process [wang2023internimage, chen2022vision, su2023towards, wang2022image, fang2023eva, tao2023siamese], which, while effective, entails significant additional costs for adaptation to a range of downstream applications. Techniques such as multi-task unification [radford2018improving, wang2022ofa, alayrac2022flamingo, wang2022git] aim to endow systems with a broad set of capabilities, yet they often fail to transcend the constraints of predetermined tasks, leaving a noticeable capability deficit in open-ended tasks when juxtaposed with LLMs. The advent of visual prompt tuning [jia2022visual, yao2021cpt, zang2022unified, wang2023images] provides a novel method for delineating specific vision tasks, like object detection, instance segmentation, and pose estimation, through the use of visual masking. Nonetheless, there is currently no work that combines LLMs with 3D lane line detection. As LLMs become more widespread and their capabilities continue to improve, research into LLM-based lane line detection presents an interesting and promising avenue for future exploration.
@@ V-B6 Towards More Accurate and Efficient 3D Lane Detection Methods for Autonomous Driving
The qualitative performance of current 3D lane detection techniques is shown in Table III. We can observe that only a few methods manage to strike a balance between model accuracy and inference latency. The experimental results of these methods indicate the need for further improvements to alleviate computational burdens while maintaining their unparalleled performance. Additionally, we select some challenging data from the 3D lane detection dataset and test the performance of the 3D lane line detection algorithm on these challenging data samples. However, the performance of the algorithm is not satisfactory, as depicted in Fig. 10. This also answers the question we raised in section I-B. Improving the algorithm’s detection performance in extreme weather conditions is also crucial. Moreover, the reported time complexities in Table III,IV and V suggest that some methods can achieve real-time execution when deployed on GPU devices. Nevertheless, considering the constrained computational resources in AD systems, the focus of 3D lane detection methods should also shift towards computational complexity.